# Collaborative Learning in Agentic Systems: A Collective AI is Greater Than the Sum of Its Parts

### v1

@@ Appendix E Implementation Details
1:Agent’s IP/Port $\texttt{addr}_{i}$, Communication graph $\mathcal{G}=\{\mathcal{I},\mathcal{E}\}$, where =$\mathcal{E}=\{(i,j)\mid i,j\in\mathcal{I},\ i\neq j\}$
@@
3:     Initialize parallel process CommunicationServer($Q_{\text{mask}}$)
@@
[…]
@@
2:     Compute embedding $v_{\tau}^{\text{new}}\leftarrow\psi(\mathcal{D})$triangleright via Eq. 6
3:     $v_{\tau}\leftarrow\frac{1}{2}(v_{\tau}+v_{\tau}^{\text{new}})$triangleright Moving average update
@@
Tables 4, 5, and 6 report all hyperparameters used for the PPO model, communication module, and embedding module, respectively. Figure 21 reports details on the model architectures used in our experiments. All seeds used in our experiments can be found in the code implementation: https://github.com/DMIU-ShELL/MOSAIC. Table 3 lists the hardware devices used in the experiments.
Figure 21: Details on architectures used for our experiments. (Left) shows the fully linear model used in the image sequence learning and MiniGrid benchmarks. (Right) shows the convolutional model used for the MiniHack benchmark. All models use ReLU activation. Blue depicts pass-through layers (these are not masked), green depicts linear masking layers, and pink depicts convolutional masking layers.
System 1
CPU AMD EPYC 9634 (168 logical cores)
GPU(s) 2× NVIDIA A100 80GB PCIe
GPU Driver 535.216.03
CUDA Version 12.2
RAM 512 GB
System 2
CPU AMD EPYC 7713P (128 logical cores)
GPU(s) 2× NVIDIA A100 40GB PCIe
GPU Driver 535.216.03
CUDA Version 12.2
RAM 256 GB
System 3
CPU AMD Ryzen 9 5950X (32 logical cores)
GPU(s) 1× NVIDIA GeForce RTX 4090
GPU Driver 535.183.01
CUDA Version 12.2
RAM 128 GB
Table 4: PPO and model hyperparameters used in our experiments across each benchmark.
Hyperparameter Image Sequence Learning MiniGrid MiniHack
Learning rate 0.00025 0.00025 0.00025
Backbone seed 9157 9157 9157
Num. workers 1 1 1
Discount factor 0.99 0.99 0.99
GAE Tau 0.99 0.99 0.99
Entropy 0.01 0.01 0.01
Rollout length 512 2048 2048
Opt. iterations 8 8 8
Num. mini batches 64 64 64
PPO ratio clip 0.1 0.1 0.1
Gradient clip 5 5 5
Total training steps 102400 1024000 2048000
Table 5: Hyperparameters specific to the communication module of the system. Querying frequency refers to the frequency of queries with respect to which iteration the agent is on. Query wait and Mask wait refer to the built-in wait delays in the peer-to-peer communication protocol. One of the primary goals of MOSAIC is asynchronicity, which means agents cannot afford to wait for extended periods for information to arrive. We found that 0.3 ms worked well in our experiments, though this may not work well when scaled to more agents.
Hyperparameter Image Sequence Learning MiniGrid MiniHack
Query freq. (iterations) 10 25 10
Query wait (ms) 0.3 0.3 0.3
Mask wait (ms) 0.3 0.3 0.3
Table 6: Hyperparameters specific to the embedding module, which generates the Wasserstein task embeddings online.
Hyperparameter Image Sequence Learning MiniGrid MiniHack
Reference size ($M$) 50 50 50
Num. samples ($N$) 128 128 128
Frequency (iterations) 1 1 1

### v2

@@ Appendix E Implementation Details
1:Agent’s IP/Port $\texttt{addr}_{i}$, Communication graph $\mathcal{G}=\{\mathcal{I},\mathcal{E}\}$, where $\mathcal{E}=\{(i,j)\mid i,j\in\mathcal{I},\ i\neq j\}$
@@
[…]
@@
Algorithm 4 Forward pass using masks for network layer $l$
2:procedure ForwardPass(x)
3:  if $|P|=0$ then $\triangleright$ no peer masks available
4:   Set $\phi_{\tau}^{\text{lc}}\leftarrow\phi_{\tau}$
5:  else
6:   $\phi_{\tau}^{\text{lc}}\leftarrow g\left(\beta_{\tau}\phi_{\tau}+\sum_{k\in P}\beta_{k}\phi_{k}\right)$ $\triangleright$ Eq. 7
7:  end if
10:  return $y$
11:end procedure
@@
6:   Accept incoming connection $(\texttt{conn},\texttt{remote\_addr})$
@@
2:  Compute embedding $v_{\tau}^{\text{new}}\leftarrow\psi(\mathcal{D})$ $\triangleright$ via Eq. 5
3:  $v_{\tau}\leftarrow\frac{1}{2}(v_{\tau}+v_{\tau}^{\text{new}})$ $\triangleright$ Moving average update